# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v1

@@ Abstract
The wide range of sub- and super-nuclear densities achieved in neutron stars makes them ideal probes of dense nuclear behavior in the form of the nuclear equation of state (EoS). Studying neutron stars both in isolation and in highly dynamic events, many recent observations, most famously the gravitational wave and electromagnetic signals associated with the BNS merger GW170817/AT2017gfo, have provided suggestive insight into these highest nuclear densities. Measurements of galactic neutron star masses and radii from NICER and other radio and X-ray measurements provide critical complementary perspectives, bounding other features of the EoS. Though nominally congruent, in this paper we highlight many underappreciated “hidden” priors embedded in joint analysis of these many messengers, and their systematic impact on joint EoS inference. In this work, we perform a careful step-by-step Bayesian inference using a simple low-dimensional parametric EoS model, incrementally adding information from galactic pulsars, gravitational wave sources, and more speculative constraints involving kilonovae and the neutron star maximum mass. At each stage, we carefully discuss the marginal likelihood, explaining how hidden priors impact conclusions. Conversely, we also quantify how much information these measurements have, arguing that many if not most have minimal impact relative to the explicit and hidden prior assumptions. Specifically, we find two features dominate our inference: on the one hand, the choice of EoS parameterization and particularly hyperprior, and on the other the astrophysical priors associated with interpreting events, particularly the multimessenger source GW170817. In an appendix, we outline a simple semianalytic projection suitable for assessing the measurability of the EoS with ensembles of present and future detections.
@@ I Introduction
Unlike many recent studies which rely on increasingly complex models for the nuclear EoS, to preserve our emphasis on the observational ingredients and priors we employ an extremely simple low-dimensional model for the nuclear EoS: two conventional spectral representations [76, 77, 78] based on approximating a functional versus (log) pressure with a Taylor series. These low-dimensional smooth approximations are known to have limited flexibility compared with more expressive infinite-dimensional nonparametric models like Gaussian processes [21, 5, 79]. However, these models are more than sufficient to express contemporary observational uncertainties, helping us indicate avenues where systematics may play an important role. Other studies have thoroughly explored more flexible approaches including multiple phase transitions; see, e..g, [74] and references therein.
This paper is organized as follows. In Section II we review our phenomenological spectral approach to the nuclear equation of state; pertinent observations of neutron stars; our specific implementation of hierarchical Bayesian inference for the nuclear equation of state; and our specific procedure and priors for each class of measurement. For galactic neutron stars we employ radio observations of heavy neutron stars PSR-J0348 [80] and PSR-J1641 [81, 82], NICER X-ray pulse profiling for PSR-J0030, PSR-J0740, and X-ray spectra and distance analyzed for HESS J1731 [83] which retains the possibility of consisting of strange quark matter, however for this study we assume it to be nucleonic. For gravitational wave sources, we employ only GW170817, emphasizing the role of hidden astrophysical priors on EoS inference. In Section III we provide inferences for the nuclear EoS using galactic pulsars and nuclear symmetry energy; update that inference using information from GW170817; and then reassess those results by incorporating more indirect interpretations of observations, including limits on the NS maximum mass. Finally, we present our concluding remarks in Sec. IV.
@@ II.1 Nuclear equation of state phenomenology
Following previous investigations, we characterise the nuclear equation of state with a low-dimensional phenomenological model. Rather than employ a sophisticated but high-dimensional phenomenological approach based on microphysics-motivated calculations with many tunable parameters such as relativistic mean field theory [84, 85, 86, 87, 88, 89, 90, 91], for simplicity and lacking observationally-driven necessity, we adopt two concrete analytic representations.
@@
For this model, we adopt a uniform prior in its hyperparameters, with limits consistent with previous studies [92, 2]. Though we do not explicitly enforce other customary constraints (e.g., on the adiabatic index), our TOV solver requires EoS that are nearly causal, so our calculations de facto reject highly unphysical samples. To illustrate the impact of model and prior systematics, we also employ another spectral formulation, referred to as $\Upsilon$-spectral from here onwards (though denoted the “improved causal spectral” formulation in original paper), designed to ensure smooth transition from the fixed low density piece to the fit higher density part. In particular we use the pressure based formulation described in [78], where the spectral expansion is constructed for the sound speed dependent parameter $\Upsilon(p)$ defined as
@@
This decomposition is very similar to the previous $\Gamma$-spectral representation [77] with the only difference being that in this new decomposition, the zeroth coefficient, $\lambda_{0}$, is predetermined. Here $\lambda_{0}$ is fixed such that sound speed continuity is ensured at the matching point, $p_{0}$, between the low density and high density pieces
@@
We restrict our analyses to the first four independent parameter in this decomposition, $k=1,2,3,4$, in order to capture systematic features of the EoSs while keeping the exploration manageable. We choose a uniform prior for these four spectral parameters. The ranges of these uniform parameter selection are $\lambda_{1}\in[-3,3.2]$, $\lambda_{2}\in[-4,2]$, $\lambda_{3}\in[-0.2,0.8]$, and $\lambda_{4}\in[-0.05,0.02]$. Figure 1 shows our implied prior on $p(\rho)$ and nonrotating neutron star gravitational mass and radius. We use these two parametric approaches to qualitatively assess the kinds of systematic uncertainty exhibited among some of the commonly used parameterizations, such as piecewise polytropic [93], dynamical-piecewise polytropic [94, 92], and similar spectral methods [76, 77, 78] which adopt different basis functions or parameters [95].
@@
We follow previous work that employs Gaussian quadrature to evaluate the energy-density at given pressure on the lalsuite implementation [2, 92]. We adopt a fixed low-density equation of state with the first four terms in the expansion determining the EoS. This parameterization has been previously constrained by comparison of its input GW170187 [2] and NICER measurements [37, 45]. For all EoSs, we solve the Tolman-Oppenheimer-Volkoff (TOV) equation [98, 99] for nonrotating neutron star structure in hydrostatic equilibrium using RePrimAnd [100, 101] to find one-parameter families of gravitational mass $M$, radius $R$, and tidal deformability $\Lambda$. The output for an equation of state $\alpha$ is characterized by a one-parameter family $R_{\alpha}(s),M_{\alpha}(s)$, where $s$ is a fiducial parameter characterizing our dimensionless prior. For simplicity, in this work we adopt the pseudo-enthalpy $h$, and only investigate the lowest stable branch of solutions.
@@
Bayesian inference frameworks provide a likelihood and hence (with a prior) probability for each equation of state function. We can therefore interpret observations as providing information about a credible interval on the space of functions: for example, deducing a volume of EoS which contain 90% of all (joint) posterior probability. As an example, Figure 1 renders our 100 percentile prior EoS distribution. This interpretation, however, conflicts with the customary approach to rendering and reporting conclusions about the EoS, which are presented as marginal distributions for (uncorrelated) point measurements: for example, the distribution of pressure at a given density, or the radius of NS of a fiducial mass. As highlighted in previous work [5], the superposition of point measurements neglects important information including correlations between different densities (for $p(\rho)$) or masses (for $R(M)$). As a concrete example for any two densities $\rho_{i}$, the posterior pressures $\ln p(\rho_{1}),\ln p(\rho_{2})$ will be correlated by measurements, above and beyond any correlations introduced by our prior. Nonetheless, in this work to simplify comparison with previously-reported results, except for Figure 1 we will typically render our conclusions using this point-measurement approach. Thus for example a confidence interval on $p(\rho)$ follows by taking all weighted EoS realizations $(p_{k}(\rho),w_{k})$ and, for each reference density $\rho_{i}$, constructing a confidence interval for $p(\rho_{i})$ using the weighted sample at that density. Similarly but not identical are confidence intervals on $\rho(p)$, constructed similarly but using the inverse relationship.
The interpretations described above assume only a prior on the functional space of EoS. However, many discussions of the NS mass-radius distribution adopt priors along the family of stable NS, by for example adopting a prior on the NS central density; see, e.g., [102, 74, 45]. As discussed below, these analyses are in effect adopting a prior on the NS mass distribution. As a result, in order to replicate their results, we must present our results in a comparable fashion. Specifically, rather than construct a 90% credible interval for the NS radius at each fixed mass, we must instead construct a two-dimensional marginal distribution in both mass and radius. This two-dimensional posterior distribution will explicitly depend on the choice of NS mass distribution (or equivalently NS central density distribution).
@@ II.2 Neutron star population models
Deducing the equation of state from observations of neutron star properties always requires some model for the neutron star population: empirical observables limit but do not uniquely determine the neutron star mass, spin, radius, and tidal deformability. In this work, unless otherwise noted, we will assume the isolated (and binary) neutron star population is uniformly distributed in gravitational mass $m$ between some fixed, EoS-independent limits $m\in[m_{min},m_{max}]$. We specifically adopt $m_{max}=2.1M_{\odot}$ and $m_{min}=0.4M_{\odot}$, with the lower and upper mass limits chosen because of the collection of NS we will use to constrain the nuclear EoS. Similarly, unless otherwise noted, we will assume the (aligned component) of neutron star spin is distributed isotropically in a sphere and uniformly in magnitude, with a maximum dimensionless value $a=S^{2}/m<0.05$. This fiducial population closely resembles the prior assumptions used when inferring the nature of low-mass gravitational wave sources. We emphasize that our prior represents our assumptions about the astrophysical distribution of neutron stars as realized in nature, which need not reach the maximum mass allowed by the nuclear equation of state.
@@
For context, other groups have adopted a wide range of different implicit or explicit priors for the neutron star mass distribution. For example, two groups adopt priors on the neutron star central density, not the neutron star mass (e.g., a quadratic prior [74, 45] or a log-uniform prior [102] between the maximum central density and some reference density). Alternatively, even within the context of nominally uniform mass distributions, other groups [103] have specifically assumed $p(m|{\cal Y})$ is uniform between a fixed minimum mass and the EoS-dependent maximum mass $M_{max}({\cal Y})$, where ${\cal Y}$ here and henceforth denotes the EoS or its parameterization. In other words, this model asserts all neutron stars that can exist do exist. In terms of the unit step function $\Theta_{I}(x)=1$ for $x\in I$ ($I$ for interval) and 0 otherwise, their one- and two-dimensional priors are $p(m)=\Theta_{I}(m)/|I|$ and $p(m_{1},m_{2})=2\Theta_{I}(m_{1})\Theta_{I}(m_{2})\Theta(m_{1}-m_{2})/|I|^{2}$, where $I=[m_{min},M_{max}({\cal Y})]$ is the interval between $m_{min}$ and the EoS-dependent maximum mass. Compared to our fiducial approach, both of these aggressive population models introduce EoS-dependent normalization factors, which weakly penalizes models with large $M_{\rm max}({\cal Y})$ for single observations and strongly penalize them for many observations. Additionally, these aggressive population models should be strongly constrained by the absence of massive neutron stars; see, e.g., Alsing et al. [104, 105] and references therein. For broader context as regards prior choices, Wysocki et al. [106] demonstrated that, if not known a priori, the neutron star mass distribution and EoS must be jointly inferred, to avoid biasing the deduced EoS. In the text and appendicies, we will briefly discuss how alternative mass priors can impacts our results.
@@
Several sources of external information provide good reason to adopt alternative informed priors. For example, several authors have proposed adopting comparable masses for binary neutron stars, motivated by galactic observations and conventional field formation scenarios which strongly favor nearly equal masses [107, 108, 109, 110, 111, 107, 108]. Similar arguments have been employed to rule out large or preferentially negatively-oriented spins [112]. (Of course, the limited statistics available from the GW census are consistent with a much broader mass ratio distribution [79, 113] and provide minimal constraints on plausible NS spin orientations; too, non-merging galactic BNS binaries have a much broader mass ratio distribution [114].) At the other extreme, because the kilonova AT2017gfo associated with GW170817 suggests very massive dynamical ejecta and since most predictions for this ejecta are suppressed for $q\simeq 1$, other authors [115, 116] have suggested favoring asymmetric masses where the dynamical ejecta isn’t suppressed. On similar grounds based on AT2017gfo, other authors [117, 118] have demonstrated that a strongly-spinning secondary neutron star ($m_{2}$) will produce substantially enhanced ejecta, as its outer layers are less bound. Strongly spinning NS mergers can occur for example if the secondary does not have time to spin down, due to fortuitous NS kicks [119] or Kozai-Lidov effects in hierarchical triples [120]. [Such a rapid formation scenario isn’t consistent with the large geometric offset for AT2017gfo but could occur in general; as noted in [120, 121] this rapid formation scenario would reconcile NS merger formation with r-process-enriched but metal-poor stars.] Particularly given the systematics associated with contemporary eject calculations, an exhaustive quantitative investigation of these additional speculative sources of prior information merits its own dedicated study.
@@
Motivated by the above discussion, we will analyze GW170817 with combinations of several alternative priors. To focus on high mass ratios only, we will perform some analyses with a mass ratio distribution $p(q)$ for AT2017gfo that is consistent with a uniform distribution of component masses but excludes $m_{2}/m_{1}>0.6$. [At GW170817’s source-frame chirp mass ${\cal M}_{c}=1.186M_{\odot}$, this constraint requires more extreme choices for source-frame primary and secondary mass: $m_{2}<1.062M_{\odot}$ and $m_{1}>1.77M_{\odot}$.] Conversely, when requiring GW170817’s mass ratio prior distribution be constrained to the range expected for GW-detected BNS sources derived from the galactic population, we require $m_{2}/m_{1}>0.9$ [107]. To allow for extreme spins, we consider an isotropic uniform spin magnitude prior which extends to $0.4$. To strongly disfavor extreme spin-orbit misalignment, we consider $\chi_{i,z}>0$. Operationally, all of these astrophysical prior choices can be implemented as alternative options for the upper and lower bounds adopted for $q,\chi_{i,z}$, when analyzing GW170817 with a nonprecessing GW model.

### v2

@@ Abstract
The wide range of sub- and super-nuclear densities achieved in neutron stars makes them ideal probes of dense nuclear behavior in the form of the nuclear equation of state (EoS). Studying neutron stars both in isolation and in highly dynamic events, many recent observations, most famously the gravitational wave and electromagnetic signals associated with the BNS merger GW170817/AT2017gfo, have provided suggestive insight into these highest nuclear densities. Measurements of galactic neutron star masses and radii from NICER and other radio and X-ray measurements provide critical complementary perspectives, bounding other features of the EoS. Though nominally congruent, in this paper we highlight many underappreciated “hidden” priors embedded in joint analysis of these many messengers, and their systematic impact on joint EoS inference. In this work, we perform a careful step-by-step Bayesian inference using a simple low-dimensional parametric EoS model, incrementally adding information from galactic pulsars, gravitational wave sources, and more speculative constraints involving kilonovae and the neutron star maximum mass. At each stage, we carefully discuss the marginal likelihood, explaining how hidden priors impact conclusions. Conversely, we also quantify how much information these measurements have, arguing that many if not most have minimal impact relative to the explicit and hidden prior assumptions. Specifically, we find two features dominate our inference: on the one hand, the choice of EoS parameterization, and on the other the astrophysical priors associated with interpreting events, particularly the multimessenger source GW170817. Spectral parameterizations have been shown to cause large differences in the predicted EoS based on if causality is applied in the framework or not, and applying causality predicts a strong phase-transition-like feature at $\sim 10^{14.5}$ g cm-3. In an appendix, we outline a simple semianalytic projection suitable for assessing the measurability of the EoS with ensembles of present and future detections.
@@ I Introduction
Unlike many recent studies which rely on increasingly complex models for the nuclear EoS, to preserve our emphasis on the observational ingredients and priors we employ an extremely simple low-dimensional model for the nuclear EoS:
two conventional spectral representations [76, 77, 78] based on approximating a functional versus (log) pressure with a Taylor series. These low-dimensional smooth approximations are known to have limited flexibility compared with more expressive infinite-dimensional nonparametric models like Gaussian processes [21, 5, 79]. However, these models are more than sufficient to express contemporary observational uncertainties, helping us indicate avenues where systematics may play an important role. Other studies have thoroughly explored more flexible approaches including multiple phase transitions; see, e..g, [74] and references therein.
This paper is organized as follows. In Section II we review our phenomenological spectral approach to the nuclear equation of state; pertinent observations of neutron stars; our specific implementation of hierarchical Bayesian inference for the nuclear equation of state; and our specific procedure and priors for each class of measurement. For galactic neutron stars we employ radio observations of heavy neutron stars PSR-J0348 [80] and PSR-J1641 [81, 82], NICER X-ray pulse profiling for PSR-J0030, PSR-J0740, and X-ray spectra and distance analyzed for HESS J1731 [83] which retains the possibility of consisting of strange quark matter, however for this study we assume it to be nucleonic. For gravitational wave sources, we employ only GW170817. In Section III we provide inferences for the nuclear EoS using galactic pulsars and nuclear symmetry energy; update that inference using information from GW170817, and then reassess those results by incorporating more indirect interpretations of observations, including limits on the NS maximum mass. Finally, we present our concluding remarks in Sec. IV.
@@ II.1 Nuclear equation of state phenomenology
Following previous investigations, we characterise the nuclear equation of state with a low-dimensional phenomenological model. Rather than employ a sophisticated but high-dimensional phenomenological approach based on microphysics-motivated calculations with many tunable parameters such as relativistic mean field theory [84, 85, 86, 87, 88, 89, 90, 91], for simplicity and lacking observationally-driven necessity, we adopt two distinct analytic representations.
@@
For this model, we adopt a uniform prior in its hyperparameters, with limits consistent with previous studies [92, 2]. Though we do not explicitly enforce other customary constraints (e.g., on the adiabatic index), our TOV solver requires EoS that are nearly causal, so our calculations de facto reject highly unphysical samples as implausible. To illustrate the impact of model and prior systematics, we also employ another spectral formulation, referred to as $\Upsilon$-spectral from here onwards (though denoted the “improved causal spectral” formulation in Ref. [78]), designed to ensure smooth transition from the fixed low density piece to the fit higher density part and to ensure causality all throughout. In particular we use the pressure based formulation described in [78], where the spectral expansion is constructed for the sound speed dependent parameter $\Upsilon(p)$ defined as
@@
This decomposition is very similar to the previous causal-$\Gamma$-spectral representation [77] with the only difference being that the zeroth coefficient, $\lambda_{0}$, is predetermined. Here $\lambda_{0}$ is fixed such that sound speed continuity is ensured at the matching point, $p_{0}$, between the low density and high density pieces
@@
We restrict our analyses to the first four independent parameter in this decomposition, $k=1,2,3,4$, in order to capture systematic features of the EoSs while keeping the exploration manageable. We choose a uniform prior for these four spectral parameters, over $\lambda_{1}\in[-3,3.2]$, $\lambda_{2}\in[-4,2]$, $\lambda_{3}\in[-0.2,0.8]$, and $\lambda_{4}\in[-0.05,0.02]$, where these ranges of each parameter are motivated by covering existing EOS as well as the $\Gamma$-spectral posterior for GW170817 [93]. Figure 1 shows our implied prior on $p(\rho)$ and nonrotating neutron star gravitational mass and radius. We use these two parametric approaches to qualitatively assess the kinds of systematic uncertainty exhibited among some of the commonly used parameterizations, such as piecewise polytropic [94], dynamical-piecewise polytropic [95, 92], and similar spectral methods [76, 77, 78] which adopt different basis functions or parameters [96].
@@
We follow previous work that employs Gaussian quadrature to evaluate the energy-density at given pressure on the lalsuite implementation [2, 92]. We adopt a fixed low-density equation of state with the first four terms in the expansion determining the EoS. This parameterization has been previously constrained by comparison of its input GW170817 [2] and NICER measurements [37, 45]. For all EoSs, we solve the Tolman-Oppenheimer-Volkoff (TOV) equation [99, 100] for nonrotating neutron star structure in hydrostatic equilibrium using RePrimAnd [101, 102] to find one-parameter families of gravitational mass $M$, radius $R$, and tidal deformability $\Lambda$. The output for an equation of state $\alpha$ is characterized by a one-parameter family $R_{\alpha}(s),M_{\alpha}(s)$, where $s$ is a fiducial parameter characterizing our dimensionless prior. For simplicity, in this work we adopt the pseudo-enthalpy h, and only investigate the lowest stable branch of solutions.
@@
Bayesian inference frameworks provide a likelihood and hence (with a prior) probability for each equation of state function. We can therefore interpret observations as providing information about a credible interval on the space of functions: for example, deducing a volume of EoS which contain 90% of all (joint) posterior probability. Figure 1 shows our 100 percentile prior EoS distribution. This interpretation, however, conflicts with the customary approach to rendering and reporting conclusions about the EoS, which are presented as marginal distributions for (uncorrelated) point measurements: for example, the distribution of pressure at a given density, or the radius of NS of a fiducial mass. As highlighted in previous work [5], the superposition of point measurements neglects important information including correlations between different densities (for $p(\rho)$) or masses (for $R(M)$). As a concrete example for any two densities $\rho_{i}$, the posterior pressures $\ln p(\rho_{1}),\ln p(\rho_{2})$ will be correlated by measurements, above and beyond any correlations introduced by our prior. Nonetheless, in this work to simplify comparison with previously-reported results, except for Figure 1 we render our conclusions using this point-measurement approach. Thus for example a confidence interval on $p(\rho)$ follows by taking all weighted EoS realizations $(p_{k}(\rho),w_{k})$ and, for each reference density $\rho_{i}$, constructing a confidence interval for $p(\rho_{i})$ using the weighted sample at that density. Similarly but not identical are confidence intervals on $\rho(p)$, constructed similarly but using the inverse relationship.
The conclusions about the EOS posterior described in the preceding paragraph assume only a prior on the functional space of EoS. However, many discussions of the NS mass-radius distribution adopt priors along the family of stable NS by, for example, adopting a prior on the NS central density; see, e.g., [103, 74, 45]. As discussed below, these analyses are in effect adopting a prior on the NS mass distribution. As a result, in order to replicate their results, we must present our results in a comparable fashion. Specifically, rather than construct a 90% credible interval for the NS radius at each fixed mass, we must instead construct a two-dimensional marginal distribution in both mass and radius. This two-dimensional posterior distribution will explicitly depend on the choice of NS mass distribution (or equivalently NS central density distribution).
@@ II.2 Neutron star population models
Deducing the equation of state from observations of neutron star properties always requires some model for the neutron star population: empirical observables limit but do not uniquely determine the neutron star mass, spin, radius, and tidal deformability. In this work, unless otherwise noted, we will assume the isolated (and binary) neutron star population is uniformly distributed in gravitational mass $m$ between fixed, EoS-independent limits $m\in[m_{min},m_{max}]$. We specifically adopt $m_{min}=0.4~{}{\rm M_{\odot}}$ and $m_{max}=2.1~{}{\rm M_{\odot}}$, with the lower and upper mass limits chosen because of the collection of NSs we will use to constrain the nuclear EoS. Similarly, unless otherwise noted, we will assume the (aligned component) of neutron star spin is distributed isotropically in a sphere and uniformly in magnitude, with a maximum dimensionless value $a=S^{2}/m<0.05$. This fiducial population closely resembles the prior assumptions used when inferring the nature of low-mass gravitational wave sources. We emphasize that our prior represents our assumptions about the astrophysical distribution of neutron stars as realized in nature, which need not reach the maximum mass allowed by the nuclear equation of state.
@@
For context, other groups have adopted a wide range of different implicit or explicit priors for the neutron star mass distribution. For example, some adopt priors on the neutron star central density, not the neutron star mass (e.g., a quadratic prior [74, 45] or a log-uniform prior [103] between the maximum central density and some reference density). Alternatively, even within the context of nominally uniform mass distributions, others [104] have specifically assumed $p(m|{\cal Y})$ is uniform between a fixed minimum mass and the EoS-dependent maximum mass $M_{max}({\cal Y})$, where ${\cal Y}$ here and henceforth denotes the EoS or its parameterization. In other words, this model asserts all NSs that can exist under an EoS do exist. In terms of the unit step function $\Theta_{I}(x)=1$ for $x\in I$ ($I$ being the interval) and 0 otherwise, their one- and two-dimensional priors are $p(m)=\Theta_{I}(m)/|I|$ and $p(m_{1},m_{2})=2\Theta_{I}(m_{1})\Theta_{I}(m_{2})\Theta(m_{1}-m_{2})/|I|^{2}$, where $I=[m_{min},M_{max}({\cal Y})]$ is the interval between $m_{min}$ and the EoS-dependent maximum mass. Compared to our fiducial approach, both of these aggressive population models introduce EoS-dependent normalization factors, which weakly penalizes models with large $M_{\rm max}({\cal Y})$ for single observations and strongly penalize them for many observations. Additionally, these aggressive population models should be strongly constrained by the absence of massive neutron stars; see, e.g., Alsing et al. [105, 106] and references therein. For broader context regarding prior choices, Wysocki et al. [93] demonstrated that, if not known a priori, the neutron star mass distribution and EoS must be jointly inferred, to avoid biasing the deduced EoS. In the main text and in appendices we will briefly discuss how alternative mass priors can impacts our results.
@@
Several sources of external information provide good reason to adopt alternative informed priors. For example, several authors have proposed adopting comparable masses for binary neutron stars, motivated by galactic observations and conventional field formation scenarios which strongly favor nearly equal masses [107, 108, 109, 110, 111, 107, 108]. Similar arguments have been employed to rule out large or preferentially negatively-oriented spins [112]. (Of course, the limited statistics available from the GW census are consistent with a much broader mass ratio distribution [79, 113] and provide minimal constraints on plausible NS spin orientations; additionally, non-merging galactic BNS binaries have a much broader mass ratio distribution [114].) At the other extreme, because the kilonova AT2017gfo associated with GW170817 suggests very massive dynamical ejecta and since most predictions for this ejecta are suppressed for $q\simeq 1$, others have suggested favoring asymmetric masses where the dynamical ejecta isn’t suppressed [115, 116]. On similar grounds based on AT2017gfo, others have demonstrated that a strongly-spinning secondary neutron star ($m_{2}$) will produce substantially enhanced ejecta, as its outer layers are less bound [117, 118]. Strongly spinning NS mergers can occur for example if the secondary does not have time to spin down, due to fortuitous NS kicks [119] or Kozai-Lidov effects in hierarchical triples [120]. (Such a rapid formation scenario isn’t consistent with the large geometric offset for AT2017gfo but could occur in general; as noted in [120, 121] this rapid formation scenario would reconcile NS merger formation with r-process-enriched but metal-poor stars.) Particularly given the systematics associated with contemporary eject calculations, an exhaustive quantitative investigation of these additional speculative sources of prior information merits its own dedicated study.
@@
Motivated by the discussion in the preceding paragraphs, we will analyze GW170817 with combinations of several alternative priors. To focus on high mass ratios only, we will perform some analyses with a mass ratio distribution $p(q)$ for AT2017gfo that is consistent with a uniform distribution of component masses but excludes $m_{2}/m_{1}>0.6$. (At GW170817’s source-frame chirp mass ${\cal M}_{c}=1.186~{}{\rm M_{\odot}}$, this constraint requires more extreme choices for source-frame primary and secondary mass: $m_{2}<1.062~{}{\rm M_{\odot}}$ and $m_{1}>1.77~{}{\rm M_{\odot}}$.) Conversely, when requiring GW170817’s mass ratio prior distribution be constrained to the range expected for GW-detected BNS sources derived from the galactic population, we require $m_{2}/m_{1}>0.9$ [107]. To allow for extreme spins, we consider an isotropic uniform spin magnitude prior which extends to $0.4$. To strongly disfavor extreme spin-orbit misalignment, we consider $\chi_{i,z}>0$. Operationally, all of these astrophysical prior choices can be implemented as alternative options for the upper and lower bounds adopted for $q,\chi_{i,z}$, when analyzing GW170817 with a nonprecessing GW model.